OpenStudy (anonymous):

Evaluate 15-4+(7-5)^2

OpenStudy (anonymous):

you have to use PEMDAS

OpenStudy (anonymous):

do you know how to use pemdas?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you use PEMDAS then you should get the answer

OpenStudy (anonymous):

I got 7

OpenStudy (anonymous):

for the answer

OpenStudy (anonymous):

what did you raise to the second power?

OpenStudy (anonymous):

2^2

OpenStudy (anonymous):

was it the whole parenthesis after you subtracted the 7-5

OpenStudy (anonymous):

let me check my math again cause that should be right

OpenStudy (anonymous):

no after the( )

OpenStudy (anonymous):

Use BODMAS Rule

OpenStudy (anonymous):

1 st u have to Solve the Bracket

OpenStudy (anonymous):

heres how I did it 15-4+(7-5)^2 15-4+4 15-8 7

OpenStudy (anonymous):

15-4+(7-5)^2 = 15 -4 + (2)^2 15 - 4 + 4 then add 19 - 4 = 15

OpenStudy (anonymous):

lol 4 -4 = 0 not 8

OpenStudy (anonymous):

mistake here @bella17 15-4+4

OpenStudy (anonymous):

15-4+4 = 15